# Machine Translation Models are Zero-Shot Detectors of Translation Direction

### v3

@@ 1 Introduction
In this work, we explore the unsupervised detection of translation directions purely on the basis of a neural machine translation (NMT) system’s translation probabilities. As illustrated in Figure 1, we hypothesize that , which, if it generally holds, would allow us to infer the original translation direction.
@@
We test the approach on 20 translation directions, experimenting with 3 massively multilingual NMT models to predict the translation probabilities of human translations, NMT-produced translations, LLM-generated translations, and pre-neural machine translations. We find that the approach detects the translation direction of human translations with an accuracy of 66% on average on the sentence level, and 80% for documents with 10 sentences. For the output of neural MT systems, detection accuracy is even higher, but our hypothesis that does not hold for the output of pre-neural systems.
@@ 2.1 Translation (Direction) Detection
To reduce reliance on in-domain supervision, unsupervised methods that rely on clustering and consequent cluster labelling have also been explored for the related task of translationese detection Rabinovich and Wintner (2015); Nisioi (2015). One could conceivably perform translation direction detection using similar methods, but this has the practical problem of requiring an expert for cluster labelling, and poor open-domain performance. In a multi-domain scenario, Rabinovich and Wintner (2015) observe that clustering based on features proposed by Volansky et al. (2013) results in clusters separated by domain rather than translation status. They address this by producing clusters, being the number of domains in their dataset, and labelling each. Clearly, labelling becomes more costly as the number of domains increases, which limits applicability to an open-domain scenario.
@@ 3 Methods
Given a parallel sentence pair , the main task in this work is to identify the translation direction between a language X and a language Y, and, consequently, establish which side is the original and which is the translation. This is achieved by comparing the conditional translation probability by an NMT model with the conditional translation probability by a model operating in the inverse direction. Our core assumption is that translations are assigned higher conditional probabilities than the original by NMT models, so if , we predict that is the translation, and the original, and the original translation direction is .
@@ 3.2 Detection on the Document Level
We also study translation direction detection on the level of documents, as opposed to individual sentences. We assume that the sentences in the document are aligned 1:1, so that we can apply an NMT model trained on the sentence level to all sentence pairs in the document, and then aggregate the result.
@@ 4.1 Unsupervised Setting
The models are architecturally similar, all being based on the Transformer architecture Vaswani et al. (2017), but they differ in the training data used, number of languages covered, and model size, and consequently in translation quality. The comparison allows conclusions about how sensitive our method is to translation quality – NLLB-200-1.3B yields the highest translation quality of the three (Tiedemann and de Gibert, 2023), but we also highlight differences in data balance. English has traditionally been dominant in the amount of training data, and all three models aim to reduce this dominance in different ways, for example via large-scale back-translation Sennrich et al. (2016) in M2M-100-418M and NLLB-200-1.3B. SMaLL-100 is a distilled version of the M2M-100-12B model, and samples training data uniformly across language pairs.
@@
Direction | Sents | Docs | HT | NMT | Pre-NMT |
cs→en | |||||
cs→uk | - | ||||
de→en | |||||
de→fr | - | ||||
en→cs | |||||
en→de | |||||
en→ru | |||||
en→uk | - | ||||
en→zh | - | ||||
fr→de | - | ||||
ru→en | |||||
uk→cs | - | ||||
uk→en | - | ||||
zh→en | - | ||||
Total |
We test the approach on datasets from the WMT news/general translation tasks from WMT16 Bojar et al. (2016), WMT22 Kocmi et al. (2022), and WMT23 Kocmi et al. (2023), which come annotated with document boundaries and the original language of each document. We include different years of WMT data not only to test the approach for different translation directions and translation types but also to rule out test set contamination for experiments with models, for which the test set predates the model release. The results for the individual WMT datasets are reported in Appendix F. We also experiment with a part of the FLORES-101 dataset Goyal et al. (2022) to test the approach on indirect translations, where English was the original language of both sides of the parallel text. We divide the data into subsets based on several categorizations:
@@ 5.1 Sentence-level Classification
A second result is that the translation detection works best for NMT-based translations (75.0% macro-average), second-best for HT (66.5% macro-average), and worst for pre-NMT (41.5% macro-average). The fact that performance for pre-neural systems is below chance level indicates that the NMT systems we use tend to assign low probabilities to the (often ungrammatical) outputs of pre-neural systems. In practice, one could pair our detection method with a monolingual model to identify such low-quality outputs.
@@ 5.5 Qualitative Analysis
A qualitative comparison of sources and translations, as illustrated in Table 8, reveals that factors such as normalization, simplification, word order interference, and sentence length influence the detection of translation direction. In Example 1, an English HT translates the German source fully, while the NMT omits half of the content, showing a high degree of simplification. Our method recognizes the simplified NMT version as a translation, but not the more complete HT. Example 2 demonstrates normalization to a high degree. This shows that MMT-models assign extremely low probabilities when the target is more ungrammatical than the source. The third example indicates that translations exhibiting normalization, simplification, and interference to a higher degree are more likely to be identified. In Example 4, source language interference in terms of word order and choice significantly impacts the detection; the more literal translation mirroring the source’s word order is recognized, while the more liberal translation is not. Finally, Example 5 highlights challenges with short sentences: The German phrase Mit freundlichen Grüßen is fairly standardized, while its French equivalents can vary in use and context, adding to the ambiguity and affecting the probability distribution in NMT. Hence, our approach fails to identify any of the French translations without additional context.
@@ 6 Conclusion
We proposed a novel approach to detecting the translation direction of parallel sentences, using only an off-the-shelf multilingual NMT system. Experiments on WMT data showed that our approach, without any task-specific supervision, is able to detect the translation direction of NMT-produced translations with relatively high accuracy, proving competitive to supervised baselines. Accuracy increases to 96% if the classifier is provided with at least 10 sentences per document. We also found a robust accuracy for translations by human translators. Finally, we applied our approach to a real-world forensic case and found that it supports the hypothesis that the English book is a forgery. Future work should explore whether our approach can be improved by mitigating directional bias of the NMT model used. Another open question is to what degree our approach will generalize to document-level translation and to translation with large language models.
@@ Appendix D Supervised Baseline (WMT) Result Validation Set with Joint Encoding
To mitigate potential input order bias, we reversed the input order for half of the training data. However, this strategy did not entirely eliminate the bias. Consequently, we report test results for both input directions in Table 16 to provide a comprehensive evaluation.
@@ Appendix E Supervised Baseline (Europarl) and Cross-Domain Application
In addition to the supervised system trained on WMT16 data, we train a supervised system on a subset of Europarl Koehn (2005); Ustaszewski (2019), a corpus consisting of parallel text from the proceedings of the European Parliament. We train a system for each language pair defr, csen, deen as described in Subsection 4.2. We train on 10,249 samples per direction, use a validation and test set of 1,281 further samples per direction each. Additionally, we apply the systems to the corresponding WMT test data subset as in Subsection 4.2, but with WMT22 defr instead of WMT22/23 enru, to test the systems’ cross-domain capabilities. We experiment in the same manner with several learning rates and epochs as in Appendix C and select the best performing system for each language pair based on the validation accuracy (defr: lr = dynamic, epoch = 5; csen: lr = dynamic, epoch = 5; deen: lr = 1e-05, epoch = 5).
@@ Appendix G Data Statistics
WMT16 | cs→en | 1499 | 1499 | |||
WMT16 | de→en | 1499 | 1499 | |||
WMT16 | en→cs | 1500 | 3000 | |||
WMT16 | en→de | 1500 | 4500 | |||
WMT16 | en→ru | 1500 | 3000 | |||
WMT16 | ru→en | 1498 | 1498 | |||
WMT22 | cs→en | - | ||||
WMT22 | cs→uk | - | ||||
WMT22 | de→en | - | ||||
WMT22 | de→fr | - | ||||
WMT22 | en→cs | - | ||||
WMT22 | en→de | - | ||||
WMT22 | en→ru | - | ||||
WMT22 | en→uk | - | ||||
WMT22 | en→zh | - | ||||
WMT22 | fr→de | - | ||||
WMT22 | ru→en | - | ||||
WMT22 | uk→cs | - | ||||
WMT22 | uk→en | - | ||||
WMT22 | zh→en | - | ||||
WMT23 | cs→uk | - | ||||
WMT23 | en→cs | - | ||||
WMT23 | en→ru | - | ||||
WMT23 | en→uk | - | ||||
WMT23 | en→zh | - | ||||
WMT23 | ru→en | - | ||||
WMT23 | uk→en | - | ||||
WMT23 | zh→en | - |
@@
bnhi | |
dehi | |
csuk | |
defr | |
zhfr | |
xhzu |
@@ Appendix H Sentence Length
Language Pair | 0-19 | 20-39 | 40-59 | 60-79 | 80-99 | 100-119 | 120-139 | 140-159 | 160-179 | 180-199 |
NMT en-cs | ||||||||||
NMT en-de | ||||||||||
NMT en-ru | ||||||||||
NMT en-uk | ||||||||||
NMT en-zh | ||||||||||
NMT cs-uk | ||||||||||
NMT de-fr | ||||||||||
Macro Avg. |

### v4

@@ 1 Introduction
In this work, we explore the unsupervised detection of translation directions purely on the basis of a multilingual neural machine translation (NMT) system’s translation probabilities. As illustrated in Figure 1, we hypothesize that , which, if it generally holds, would allow us to infer the original translation direction.
@@
We test the approach on 20 translation directions, experimenting with 3 massively multilingual NMT models to predict the translation probabilities of human translations, NMT-produced translations, LLM-generated translations, and pre-neural machine translations. We find that the approach detects the translation direction of human translations with an average accuracy of 66% at the sentence level, and 80% for documents with 10 sentences. For the output of NMT systems, detection accuracy is even higher, but our hypothesis that does not hold for the output of pre-neural systems.
@@ 2.1 Translation (Direction) Detection
To reduce reliance on in-domain supervision, unsupervised methods that rely on clustering and consequent cluster labelling have also been explored for the related task of translationese detection Rabinovich and Wintner (2015); Nisioi (2015). One could conceivably perform translation direction detection using similar methods, but this has the practical problem of requiring an expert for cluster labelling and poor open-domain performance. In a multi-domain scenario, Rabinovich and Wintner (2015) observe that clustering based on features proposed by Volansky et al. (2013) results in clusters separated by domain rather than translation status. They address this by producing clusters, being the number of domains in their dataset, and labelling each. Clearly, labelling becomes more costly as the number of domains increases, which limits applicability to an open-domain scenario.
@@ 3 Methods
Given a parallel sentence pair , the main task in this work is to identify the translation direction between a language X and a language Y, and, consequently, establish which side is the original and which is the translation. This is achieved by comparing the conditional translation probability by an NMT model with the conditional translation probability by a model operating in the inverse direction. Our core assumption is that segment pairs in the original translation direction are assigned higher conditional probabilities by NMT models than in the inverse direction, so if , we predict that is the translation, and the original, and the original translation direction is .
@@ 3.2 Detection on the Document Level
We also study translation direction detection at the level of documents, as opposed to individual sentences. We assume that the sentences in the document are aligned 1:1, so that we can apply an NMT model trained at the sentence level to all sentence pairs in the document, and then aggregate the result.
@@ 4.1 Unsupervised Setting
The models are architecturally similar, all being based on the Transformer architecture Vaswani et al. (2017), but they differ in the training data used, number of languages covered, model size, and, consequently, in translation quality. The comparison allows conclusions about how sensitive our method is to translation quality – NLLB-200-1.3B yields the highest translation quality of the three (Tiedemann and de Gibert, 2023), but we also highlight differences in data balance. English has traditionally been dominant in the amount of training data, and all three models aim to reduce this dominance in different ways, for example via large-scale back-translation Sennrich et al. (2016) in M2M-100-418M and NLLB-200-1.3B. SMaLL-100 is a distilled version of the M2M-100-12B model, and samples training data uniformly across language pairs.
@@
Direction | Sents | Docs (10) | HT | NMT | Pre-NMT |
csen | 1448 | 129 | 2896 | 15928 | 16489 |
csuk | 3947 | 112 | 3947 | 49381 | - |
deen | 1984 | 121 | 3968 | 17856 | 13491 |
defr | 1984 | 73 | 1984 | 11904 | - |
encs | 4111 | 204 | 6148 | 51480 | 27000 |
ende | 2037 | 125 | 4074 | 18333 | 18000 |
enru | 4111 | 174 | 4111 | 47295 | 15000 |
enuk | 4111 | 174 | 4111 | 41147 | - |
enzh | 4111 | 204 | 6148 | 57591 | - |
frde | 2006 | 71 | 2006 | 14042 | - |
ruen | 3739 | 136 | 3739 | 40836 | 13482 |
ukcs | 2812 | 43 | 2812 | 33744 | - |
uken | 3844 | 88 | 3844 | 40266 | - |
zhen | 3851 | 162 | 5726 | 52140 | - |
Total | 44096 | 1816 | 55514 | 491943 | 103462 |
We test the approach on datasets from the WMT news/general translation tasks from WMT16 Bojar et al. (2016), WMT22 Kocmi et al. (2022), and WMT23 Kocmi et al. (2023), which come annotated with document boundaries and the original language of each document. We include different years of WMT data not only to test the approach for different translation directions and translation types but also to rule out test set contamination for experiments with models, for which the test set predates the model release. The results for the individual WMT datasets are reported in Appendix F.
We also experiment with a subset of the FLORES-101 dataset Goyal et al. (2022) to test the approach on indirect translations, where English was the original language of both sides of the parallel text. We divide the data into subsets based on several categorizations:
@@ 5.1 Sentence-level Classification
A second result is that the translation detection works best for NMT-based translations (75.0% macro-average), second-best for HT (66.5% macro-average), and worst for pre-NMT (41.5% macro-average). The fact that performance for pre-neural systems is below chance level indicates that the NMT systems we use tend to assign low probabilities to the (often ungrammatical) outputs of pre-neural systems.
@@ 5.5 Qualitative Analysis
A qualitative comparison of sources and translations, as illustrated in Table 8, reveals that factors such as normalization, simplification, word order interference, and sentence length influence the detection of translation direction. In Example 1, an English HT translates the German source fully, while the NMT omits half of the content, showing a high degree of simplification. Our method recognizes the simplified NMT version as a translation, but not the more complete HT. Example 2 demonstrates normalization to a high degree. This shows that NMT-models assign extremely low probabilities when the target is more ungrammatical than the source. The third example indicates that translations exhibiting normalization, simplification, and interference to a higher degree are more likely to be identified. In Example 4, source language interference in terms of word order and choice significantly impacts the detection; the more literal translation mirroring the source’s word order is recognized, while the more liberal translation is not. Finally, Example 5 highlights challenges with short sentences: The German phrase Mit freundlichen Grüßen is fairly standardized, while its French equivalents can vary in use and context, adding to the ambiguity and affecting the probability distribution in NMT. Hence, our approach fails to identify any of the French translations without additional context.
@@ 6 Conclusion
We proposed a novel approach to detecting the translation direction of parallel texts, using only an off-the-shelf multilingual NMT system. Experiments on WMT data showed that our approach, without any task-specific supervision, is able to detect the translation direction of NMT-produced translations with relatively high accuracy, proving competitive to supervised baselines. Accuracy increases to 96% if the classifier is provided with at least 10 sentences per document. We also found a robust accuracy for translations by human translators. Finally, we applied our approach to a real-world forensic case and found that it supports the hypothesis that the English book is a forgery. Future work should explore whether our approach can be improved by mitigating directional bias of the NMT model used. Another open question is to what degree our approach will generalize to document-level translation.
@@
Acknowledgements
We thank the anonymous reviewers for their helpful comments and suggestions. This project was funded by the Swiss National Science Foundation (project MUTAMUR; no. 213976 and project InvestigaDiff; no. 10000503).
@@ Appendix D Supervised Baseline (WMT) Result Validation Set with Joint Encoding
To mitigate potential input order bias, we reversed the input order for half of the training data. However, a comparison of the results for the WMT test set with differing input order of the source and the target (Table 16) shows that this strategy did not eliminate the bias.
@@ Appendix E Supervised Baseline (Europarl) and Cross-Domain Application
In addition to the supervised system trained on WMT16 data, we train a supervised system on a subset of Europarl Koehn (2005); Ustaszewski (2019), a corpus consisting of parallel text from the proceedings of the European Parliament. We train a system for each language pair defr, csen, deen as described in Subsection 4.2. We train on 10,249 samples per direction, use a validation and test set of 1,281 further samples each per direction. Additionally, we apply the systems to the corresponding WMT test data subset as in Subsection 4.2, but with WMT22 defr instead of WMT22/23 enru, to test the systems’ cross-domain capabilities. We experiment in the same manner with several learning rates and epochs as in Appendix C and select the best performing system for each language pair based on the validation accuracy (defr: lr = dynamic, epoch = 5; csen: lr = dynamic, epoch = 5; deen: lr = 1e-05, epoch = 5).
@@ Appendix G Data Statistics
WMT16 | csen | 1499 | 40 | 1499 | 1499 | 16489 |
WMT16 | deen | 1499 | 55 | 1499 | 1499 | 13491 |
WMT16 | encs | 1500 | 54 | 1500 | 3000 | 27000 |
WMT16 | ende | 1500 | 54 | 1500 | 4500 | 18000 |
WMT16 | enru | 1500 | 54 | 1500 | 3000 | 15000 |
WMT16 | ruen | 1498 | 52 | 1498 | 1498 | 13482 |
WMT22 | csen | 1448 | 129 | 2896 | 15928 | - |
WMT22 | csuk | 1930 | 13 | 1930 | 23160 | - |
WMT22 | deen | 1984 | 121 | 3968 | 17856 | - |
WMT22 | defr | 1984 | 73 | 1984 | 11904 | - |
WMT22 | encs | 2037 | 125 | 4074 | 20370 | - |
WMT22 | ende | 2037 | 125 | 4074 | 18333 | - |
WMT22 | enru | 2037 | 95 | 2037 | 22407 | - |
WMT22 | enuk | 2037 | 95 | 2037 | 18333 | - |
WMT22 | enzh | 2037 | 125 | 4074 | 26481 | - |
WMT22 | frde | 2006 | 71 | 2006 | 14042 | - |
WMT22 | ruen | 2016 | 73 | 2016 | 20160 | - |
WMT22 | ukcs | 2812 | 43 | 2812 | 33744 | - |
WMT22 | uken | 2018 | 22 | 2018 | 20180 | - |
WMT22 | zhen | 1875 | 102 | 3750 | 22500 | - |
WMT23 | csuk | 2017 | 99 | 2017 | 26221 | - |
WMT23 | encs | 2074 | 79 | 2074 | 31110 | - |
WMT23 | enru | 2074 | 79 | 2074 | 24888 | - |
WMT23 | enuk | 2074 | 79 | 2074 | 22814 | - |
WMT23 | enzh | 2074 | 79 | 2074 | 31110 | - |
WMT23 | ruen | 1723 | 63 | 1723 | 20676 | - |
WMT23 | uken | 1826 | 66 | 1826 | 20086 | - |
WMT23 | zhen | 1976 | 60 | 1976 | 29640 | - |
@@
bnhi | 1012 |
dehi | 1012 |
csuk | 1012 |
defr | 1012 |
zhfr | 1012 |
xhzu | 1012 |
@@ Appendix H Sentence Length
Language Pair | 0–19 | 20–39 | 40–59 | 60–79 | 80–99 | 100–119 | 120–139 | 140–159 | 160–179 | 180–199 |
NMT en-cs | 59.41 | 62.46 | 65.88 | 69.62 | 74.42 | 75.28 | 74.93 | 78.92 | 79.08 | 81.45 |
NMT en-de | 63.97 | 67.20 | 66.24 | 71.99 | 76.38 | 76.21 | 79.26 | 79.34 | 80.50 | 84.60 |
NMT en-ru | 64.82 | 62.99 | 68.73 | 71.80 | 75.44 | 76.50 | 79.42 | 79.12 | 81.16 | 81.33 |
NMT en-uk | 64.59 | 67.58 | 72.44 | 75.31 | 78.63 | 79.67 | 81.06 | 78.85 | 84.22 | 79.90 |
NMT en-zh | 74.32 | 73.82 | 77.00 | 81.31 | 82.84 | 83.75 | 84.33 | 83.49 | 86.23 | 87.12 |
NMT cs-uk | 61.93 | 69.72 | 76.24 | 76.39 | 80.21 | 80.69 | 83.42 | 81.46 | 83.07 | 84.11 |
NMT de-fr | 57.44 | 65.25 | 70.20 | 74.27 | 76.20 | 73.70 | 79.93 | 72.60 | 76.90 | 78.63 |
Macro Avg. | 63.78 | 67.00 | 70.96 | 74.38 | 77.73 | 77.97 | 80.34 | 79.11 | 81.59 | 82.45 |